# In Search of Truth: An Interrogation Approach to Hallucination Detection

### v2

@@ Abstract
Despite the many advances of Large Language Models (LLMs) and their unprecedented rapid evolution, their impact and integration into every facet of our daily lives is limited due to various reasons. One critical factor hindering their widespread adoption is the occurrence of hallucinations, where LLMs invent answers that sound realistic, yet drift away from factual truth. In this paper, we present a novel method for detecting hallucinations in large language models, which tackles a critical issue in the adoption of these models in various real-world scenarios. Through extensive evaluations across multiple datasets and LLMs, including Llama-2, we study the hallucination levels of various recent LLMs and demonstrate the effectiveness of our method to automatically detect them. Notably, we observe up to 62% hallucinations for Llama-2 in a specific experiment, where our method achieves a Balanced Accuracy (B-ACC) of 87%, all without relying on external knowledge  [^1] Our code, datasets, and task prompts can be found here. .
@@ 1 Introduction
In recent years, the emergence of LLMs such as GPT-3 Brown et al. (2020b), PaLM Chowdhery et al. (2022), and LLama Touvron et al. (2023a, b) has revolutionized natural language processing. These models enable machines to understand and generate human-like text with unprecedented fluency and coherence. These models, trained on vast amounts of text data, have demonstrated remarkable capabilities in various applications, from automated content generation to virtual assistants, and beyond. However, their remarkable performance comes with a set of challenges and concerns that need to be addressed for their responsible and effective use. A major concern is the phenomenon of hallucination, whereby these language models generate misleading, potentially harmful, or erroneous text. Hallucination can be characterized by the presence of false information in the output generated by the language model that lacks a factual basis. There are significant challenges associated with the deployment of large language models in real-world applications, especially in those involving critical information or decision-making processes.
@@ 2 Related Work
To address inconsistencies in generated text, SelfCheckGPT, introduced by Manakul et al. (2023b), leverages multiple stochastic samples generated by LLMs using the same query. SelfCheckGPT evaluates the coherence between the response and the stochastic samples by querying the same LLM multiple times. Specifically, it employs an additional prompt that includes a stochastic sample and a sentence from the generated text and predicts whether the sentence is supported by the stochastic sample. The methodology of SelfCheckGPT encompasses two distinct approaches, one is based on BERTScore Fu et al. (2023), and the other employs a multiple-choice question answering and generation approach (MQAG) Manakul et al. (2023a).
In recent research, Azaria and Mitchell (2023) proposed a method employing a multi-layer perceptron classifier that uses hidden representations from language models to predict sentence truthfulness. However, this approach necessitates labeled data for supervised training and access to the internal states of the language model, which may not always be readily available.
In Kadavath et al. (2022), the authors present a self-evaluation technique where models are trained to predict their knowledge of the answer to any given free-form question. This approach entails prompting the language model to internally assess the accuracy of its previous predictions, including estimating the likelihood that its generated response or answer is correct. It is worth noting that this method requires labeled data for model training, making it a supervised task, which differs from our settings.
@@ 3 Problem setup
We assume a source domain of textual queries and a target domain of textual answers. A few-shot prompt [^2] While our approach assumes a provided few-shot prompt, it stays adaptable to many zero-shot tasks where the creation of few-shot prompts is feasible. Brown et al. (2020b), a corresponding query $Q$ and a LLM denoted by $F_{LLM}$, are provided. The query is constructed on top of the prompt and fed into the LLM to generate an answer to the query. Our task is to detect whether the generated answer suffers from hallucinations.
@@ 4 The InterrogateLLM method
1 Step 1: Query Forward Pass $\mathcal{P}_{forward}=q_{1},a_{1},\ldots,q_{n},a_{n}$  $A^{*}=F_{LLM}([\mathcal{P}_{forward};Q])$ Step 2: Query Reconstruction $\mathcal{P}_{backward}=a_{1},q_{1},\ldots,a_{n},q_{n}$  $Q^{*}=\{\}$ for $i=1$ to $K$ do
@@ 2       Substep 1: Reverse Pass $Q^{*}=Q^{*}\cup B_{LLM}([\mathcal{P}_{backward};A^{*}])$
[…]
@@
In InterrogateLLM , the backward process is repeated multiple times ($K$ times for each model in $B_{LLM}$, see Sec. 5.2 for more details). Therefore,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|Q^{*}|=K*|B_{LLM}|}$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Q^{*}_{vec}=\{M_{embedding}(q^{*}):\forall q^{*}\in Q^{*}\}}$
@@ 4.1 Variable temperatures
$F_{LLM}$
Movies
Books
GCI
GPT3
37%
38%
0%
Llama-2 (7B)
87%
66%
25%
Llama-2 (13B)
72%
58%
60%
@@
Table 2: Hallucination detection results for all models and datasets. InterrogateLLM is reported with $K=5$ and variable temperature values. For each dataset and $F_{LLM}$, we compare InterrogateLLM and its variants to all other baselines. As GPT3 does not suffer from hallucinations on the GCI dataset, only the B-ACC metric is reported.
@@ 5.1.2 Books Dataset
The second dataset ("books dataset [^4] Books Dataset)" is derived from Amazon and includes over 200,000 literary books. This public dataset provides an overview of diverse literary books available on the Amazon platform. Each record includes details like book title, authors, publishers, and publication year.
@@ 5.1.3 Global Country Information (GCI)
The “Global Country Information” [^5] GCI Dataset (GCI) is a public dataset containing information on 195 countries. Detailed information about each country is provided, including its name, land area, capital or major city, GDP, and more. This dataset offers a comprehensive representation of global country information. In the GCI dataset, we concentrate on country and capital pairs. The task involves determining a country’s capital by asking, "What is the capital of $x$?"
Samples from the above three datasets can be found in the supplementary Sec. B. The prompts used in each datasets and the reversed prompts created by InterrogateLLM can be found in the code.
@@ 5.2 Implementation details
We set $K=5$ and $\tau=0.91$ across all experiments. Maintaining a relatively small value for $K$ facilitates rapid benchmarking of various models on datasets in our evaluations, that encompass tens of thousands of generated answers. The hyperparameter $\tau$ was determined through an analysis of ada002 embeddings on a third-party dataset. This involved embedding both similar and dissimilar sentence pairs within the QQP dataset Chen et al. (2018) and selecting the optimal threshold that effectively distinguished between the two distributions. The initial temperature $T_{0}$ was set to the default temperature of each of the evaluated LLMs, specifically $0.6$ for GPT3 and both Llama-2 models. The embedding model used in InterrogateLLM leverages the latest OpenAI’s model, ada002666https://platform.openai.com/docs/guides/embeddings/use-cases.
In our experiments, we utilized a single A100 GPU. The execution time for a single application of InterrogateLLM , with the full method, leveraging an ensemble of three models, takes up to 2 seconds. Consequently, benchmarking InterrogateLLM across the three datasets takes up to $\sim 3.44$ hours. Further insights into the hyperparameters and experimental environment will be detailed in the following subsections.
@@ 5.4 The hallucination rates
We evaluate InterrogateLLM on answers generated by three recent LLMs for each of the datasets and tasks described above. The LLMs we evaluate are: GPT-3 Brown et al. (2020b) and Llama-2 Touvron et al. (2023b) (7b and 13b chat models). Interestingly, in Tab. 1 we report the hallucination rates in the generated answers of the three models across the different datasets. Notably, GPT-3 exhibits a lower hallucination rate across all datasets and tasks, compared to the Llama-2 models.
@@ 5.5 Hallucination detection results
Binary predictions (hallucinations or not) are compared to the ground truth test labels of each dataset. For each dataset and task, we employ InterrogateLLM with four different LLM choices for the backward step: GPT-3, Llama-2 (7B), and Llama-2 (13B), either individually or as ensembles of all three models. In Tab. 2, we report the area under the curve (AUC) of the receiver operating characteristic, and the balanced accuracy (B-ACC) metrics.
As can be seen in the table, the different variants of our method improve upon all the baselines by a sizeable margin. Within the context of the InterrogateLLM method’s different variants, our findings highlight GPT-3 as a superior $B_{LLM}$ compared to both Llama-2 variants across both the Movies and Books datasets. Notably, in cases where the $B_{LLM}$ performance with Llama-2 variants aligns closely with that of GPT-3, the ensemble approach proves highly advantageous. This implies that the combination of high-performing models further enhances the overall accuracy of the technique. Furthermore, when the ensemble attains performance levels similar to the GPT-3 backward model, we find that the Llama-2 variants exhibit weaker individual performance. In such instances, the ensemble effectively mitigates the weaknesses inherent in the Llama-2 models, resulting in an overall performance that closely aligns with that of the best single model. Additional insights into limitations are provided in Sec.6.
@@ 5.6 Ablation and hyper-parameter analysis
The performance of InterrogateLLM with various values of $K$ is evaluated on the Movies, Books, and GCI datasets, and the results are reported in Tab. 3, and Tab. 9, 10 from the supplementary, respectively. Specifically, in this study, we evaluate InterrogateLLM with $K$ taking values in the range $[1,...,5]$ (higher $K$ values can be considered at the expense of more compute power). The tables reveal that utilizing $K>1$ in the backward step is crucial in all three experiments. Notably, the best results are consistently obtained with higher $K$ value. Therefore, we hypothesize that increasing the value of $K$ could potentially enhance the results, albeit at the expense of additional computational resources.
Fig. 2 depicts the enhancements araise from the different values of $K$, shown for each dataset seperately, and reported with both AUC and B-ACC metrics. Each data point represents the average result across all three forward LLMs along with all their corresponding backward LLMs (i.e. the average of each column in tables 3,9 and 10). As can be seen, the data reveals a consistent trend wherein the cumulative improvements exhibit a proportional relationship with the size of $K$.
@@
We extend our investigation to varying temperatures for the backward process. For each index $i\in\text{range}(K)$, the InterrogateLLM method utilizes a variable temperature $T_{i}$ as defined in Eq.(8). This adjustment in temperature aimed to augment the creativity and stochastic aspects of the backward model $B_{LLM}$ throughout the query reconstruction process, fostering the generation of a more diverse set of reconstructed queries. Tab. 4, and Tab. 6, 7 from the supplementary Sec.A.3, present the results of InterrogateLLM with $K=5$, when using the same temperature through all the backward processes versus using variable temperatures, as proposed in InterrogateLLM . As can be seen, the variable temperature improves the results across most experiments in the Movies dataests, while yielding on par performance in the Books and GCI datasets (see Tab.6, 7). We hypothesize that the introduction of variable temperature, generating reconstructions with diverse levels of creativity, can be particularly helpful in mitigating instances of mode "collapse", in which certain backward models consistently generate identical reconstructions. In such cases, the incorporation of variable temperature becomes more important. The proposed method utilizes a diverse range of reconstructions. When the vast majority of these diverse reconstructions closely align with the original query, it signifies robust backward processes, better reflecting a non-hallucinated answer, and consequently leading to improved accuracy scores. More ablative experiments related to the choices made in Eq. 5 can be found in the sup. Sec.A.1.
@@ 6 Limitations
7 Conclusion
As a future work, we would like to extend the method to Retrieval Augmented Generation settings, where a query is provided with a retrieved-context, and the task is to generate an answer based on the provided information in the context.
@@ A.1 The Average-Max Analysis
For each query, we obtain a list of size K containing cosine similarity scores, representing the similarity between the generated query and the original query. To ascertain the closeness of the generated query to the original, we employ two strategies: the maximum (Max) and the average (Average) of the K cosine similarity scores. Notably, in the majority of cases, the average outperformed the maximum, indicating a more robust measure for query similarity. We attribute this observation to the sensitivity of the maximum strategy to noisy cosine similarity scores, which tend to be prevalent in certain instances. Our analysis suggests that the average is a more reliable indicator of query similarity, providing a stable and effective metric for the evaluation of our reconstruction approach. Full results with avg/max are presented in tables 8,9 and 10.

### v3

# InterrogateLLM: Zero-Resource Hallucination Detection in LLM-Generated Answers

@@ Abstract
Despite the many advances of Large Language Models (LLMs) and their unprecedented rapid evolution, their impact and integration into every facet of our daily lives is limited due to various reasons. One critical factor hindering their widespread adoption is the occurrence of hallucinations, where LLMs invent answers that sound realistic, yet drift away from factual truth. In this paper, we present a novel method for detecting hallucinations in large language models, which tackles a critical issue in the adoption of these models in various real-world scenarios. Through extensive evaluations across multiple datasets and LLMs, including Llama-2, we study the hallucination levels of various recent LLMs and demonstrate the effectiveness of our method to automatically detect them. Notably, we observe up to 87% hallucinations for Llama-2 in a specific experiment, where our method achieves a Balanced Accuracy of 81%, all without relying on external knowledge  [^1] Our code, datasets, and task prompts can be found here. .
@@ 1 Introduction
In recent years, the emergence of LLMs such as GPT-3 Brown et al. (2020), PaLM Chowdhery et al. (2022), and LLama Touvron et al. (2023a, b) has revolutionized natural language processing. These models enable machines to understand and generate human-like text with unprecedented fluency and coherence. Trained on vast amounts of text data, they have demonstrated remarkable capabilities in various applications, from automated content generation to virtual assistants, and beyond. However, their remarkable performance comes with a set of challenges and concerns that need to be addressed for their responsible and effective use. A major concern is the phenomenon of hallucination, whereby these language models generate misleading, potentially harmful, or erroneous text. Hallucination can be characterized by the presence of false information in the output generated by the language model that lacks a factual basis. There are significant challenges associated with the deployment of large language models in real-world applications, especially in those involving critical information or decision-making processes.
@@ 2 Related Work
To address inconsistencies in generated text, SelfCheckGPT, introduced by Manakul et al. (2023b), leverages multiple stochastic samples generated by LLMs using the same query. SelfCheckGPT evaluates the coherence between the response and the stochastic samples by querying the same LLM multiple times. Specifically, it incorporates an additional prompt that includes a stochastic sample and a sentence from the generated text and predicts whether the sentence is supported by the stochastic sample. The approach validates each sentence by conditioning the LLM on each stochastic sample. The methodology of SelfCheckGPT encompasses various approaches, including one based on BERTScore Fu et al. (2023), and another employing a multiple-choice question answering and generation approach (MQAG) Manakul et al. (2023a), as well as n-gram and LLM-Prompting. Our method is benchmarked against this baseline, using the last approach in our study.
In recent research, Azaria and Mitchell (2023) proposed a method employing a multilayer perceptron classifier that uses hidden representations from language models to predict sentence truthfulness. However, this approach necessitates labeled data for supervised training and access to the internal states of the language model, which may not always be readily available. In Kadavath et al. (2022), the authors present a self-evaluation technique where models are trained to predict their knowledge of the answer to any given free-form question. This approach entails prompting the language model to internally assess the accuracy of its previous predictions, including estimating the likelihood that its generated response or answer is correct. It is worth noting that this method requires labeled data for model training, making it a supervised task, which differs from our settings.
@@ 3 Problem setup
We assume a source domain of textual queries and a target domain of textual answers. A few-shot prompt [^2] While our approach assumes a provided few-shot prompt, it stays adaptable to many zero-shot tasks where the creation of few-shot prompts is feasible. Brown et al. (2020), a corresponding query $Q$ and a LLM denoted by $F_{LLM}$, are provided. The query is constructed on top of the prompt and fed into the LLM to generate an answer to the query. Our task is to detect whether the generated answer suffers from hallucinations.
@@ 4 The InterrogateLLM method
In InterrogateLLM, this backward process is repeated multiple times ($K$ times for each model in $B_{LLM}$, see Sec. 5.2 for more details), with variable temperature values, as explained below. Therefore,
$|Q^{*}|=K*|B_{LLM}|$
@@
1
2Step 1: Query Forward Pass $\mathcal{P}_{forward}=q_{1},a_{1},\ldots,q_{n},a_{n}$  $A^{*}=F_{LLM}([\mathcal{P}_{forward};Q])$
3Step 2: Query Reconstruction $\mathcal{P}_{backward}=a_{1},q_{1},\ldots,a_{n},q_{n}$
[…]
5for $i=1$ to $K$ do
@@ 6       Substep 1: Reverse Pass $Q^{*}=Q^{*}\cup B_{LLM}([\mathcal{P}_{backward};A^{*}])$
7Step 3: Text Embedding $Q_{vec}=M_{embedding}(Q)$ $Q^{*}_{vec}=\{M_{embedding}(q^{*}):\forall q^{*}\in Q^{*}\}$
8Step 4: Verification $\text{sim}(Q,Q^{*})=\underset{\forall q^{*}_{\text{vec}}\in Q^{*}_{\text{vec}}}{\text{AVG}}\left(\left[C(Q_{\text{vec}},q^{*}_{\text{vec}})\right]\right)$
9if $\text{sim}(Q,Q^{*})\geq\tau$ then
@@
$Q^{*}_{vec}=\{M_{embedding}(q^{*}):\forall q^{*}\in Q^{*}\}$
@@ 4.1 Variable temperatures
$F_{LLM}$ Movies Books GCI
GPT3 37% 38% 0%
Llama-2 (7B) 87% 66% 25%
Llama-2 (13B) 72% 58% 60%
@@
SelfCheckGPT 0.794 0.689 0.751 0.693 0.934 0.891
Table 2: Hallucination detection results for all models and datasets. InterrogateLLM is reported with $K=5$ and variable temperature values. For each dataset and $F_{LLM}$, we compare InterrogateLLM and its variants to all other baselines. As GPT3 does not suffer from hallucinations on the GCI dataset, only the ACC metric is reported (in the B-ACC column).
@@ 5.1.2 Books Dataset
The second dataset ("books dataset") [^4] Books Dataset is derived from Amazon and includes over 200,000 literary books. This public dataset provides an overview of diverse literary books available on the Amazon platform. Each record includes details like book title, authors, publishers, and publication year.
@@ 5.1.3 Global Country Information (GCI)
The “Global Country Information” [^5] GCI Dataset (GCI) is a public dataset containing information on 181 countries. Detailed information about each country is provided, including its name, land area, capital or major city, GDP, and more. This dataset offers a comprehensive representation of global country information. In the GCI dataset, we concentrate on country and capital pairs. The task involves determining a country’s capital by asking, "What is the capital of $x$?"
Samples from the above three datasets can be found in the supplementary Sec. B. The prompts used in each dataset and the reversed prompts created by InterrogateLLM, can be found in the code666 GitHub project.
@@ 5.2 Implementation details
We set $K=5$ and $\tau=0.91$ across all experiments. Maintaining a relatively small value for $K$ facilitates rapid benchmarking of various models on datasets in our evaluations, that encompass tens of thousands of generated answers. The hyperparameter $\tau$ was determined through an analysis of ada002 embeddings on a third-party dataset. This involved embedding both similar and dissimilar sentence pairs within the QQP dataset Chen et al. (2018) and selecting the optimal threshold that effectively distinguished between the two distributions. The initial temperature $T_{0}$ was set to the default temperature of each of the evaluated LLMs, specifically $0.6$ for GPT3 and both Llama-2 models. The embedding model used in InterrogateLLM leverages the latest OpenAI’s model, ada002777https://platform.openai.com/docs/guides/embeddings/use-cases.
In our experiments, we used one A100 GPU. A single application of InterrogateLLM with the full method for $k=1$, using an ensemble of three models, takes up to 2 seconds. Consequently, benchmarking InterrogateLLM across the three datasets takes up to $\sim 3.44$ hours. Further insights into the hyperparameters and experimental environment will be detailed in the following subsections.
@@ 5.3 Baselines
SelfCheckGPT with Prompt: utilizes the same $F_{LLM}$ in each task, SelfCheckGPT generates additional $N$ stochastic LLM response samples, denoted as ${S_{1},S_{2},...,S_{n}}$, using the same query. Then, it scores the consistency between the generated response and the stochastic samples, by querying an LLM to determine whether the $i$-th sentence in $A^{*}$ is supported by the corresponding sample $S_{i}$. The final inconsistency score is computed by averaging the sentence scores. In the experiments, this scoring step is evaluated using GPT-3 for all tasks.
@@ 5.4 The hallucination rates
We evaluate InterrogateLLM on answers generated by three recent LLMs for each of the datasets and tasks described above. The LLMs we evaluate are: GPT-3 Brown et al. (2020) and Llama-2 Touvron et al. (2023b) (7b and 13b models). Interestingly, in Tab. 1 we report the hallucination rates in the generated answers of the three models across the different datasets. Notably, GPT-3 exhibits a lower hallucination rate across all datasets and tasks, compared to the Llama-2 models.
@@ 5.5 Hallucination detection results
Binary predictions (hallucinations or not) are compared to the ground truth test labels of each dataset.
For each dataset and task, we employ InterrogateLLM with four different LLM choices for the backward step: GPT-3, Llama-2 (7B), and Llama-2 (13B), either individually or as ensembles of all three models. In Tab. 2, we report the area under the curve (AUC) of the receiver operating characteristic and balanced accuracy (B-ACC) metrics.
As can be seen in the table, the different variants of our method improve upon all the baselines by a sizeable margin. Importantly, we note sizeable improvements also in comparison to SelfCheckGPT. This advantage attributed to InterrogateLLM stems from predicting the query back using the few-shot samples provided in the prompt, a factor entirely overlooked by SelfCheckGPT. Additionally, we observed that in many instances of hallucinations, the stochastic samples generated by SelfCheckGPT also exhibited the same mistake. Therefore, the SelfCheckGPT algorithm erroneously predicted the hallucinated $A^{*}$ as factual truth. This emphasizes the importance of our unique backward validation strategy, which differs from the query that initially caused the hallucination. Within the variants of the InterrogateLLM method, we observe that the use of an ensemble in the backward process exhibits sizeable strides across the board, suggesting that model diversity can compensate for individual model weaknesses (see also Sec.7).
@@ 5.6 Ablation and hyper-parameter analysis
The performance of InterrogateLLM with various values of $K$ is evaluated on the Movies, Books, and GCI datasets, and the results are reported in Tab. 3, and Tab. 9, 10 from the supplementary, respectively. Specifically, in this study, we evaluate InterrogateLLM with $K$ taking values in the range $[1,...,5]$ (higher $K$ values can be considered at the expense of more compute power). The tables reveal that utilizing $K>1$ in the backward step is crucial in all three experiments. Notably, the best results are consistently obtained with higher $K$ value, where $K=5$ takes the lead in the majority of cases. Therefore, we hypothesize that increasing the value of $K$ could potentially enhance the results, albeit at the expense of additional computational resources. In addition, we observe that the ensemble of all three models (GPT-3, Llama-2 (7B), and Llama-2 (13B)) yielded the highest performance across all $K$ values. This suggests once again that combining recovery scores from multiple models enhances hallucination detection.
Fig. 2 depicts the enhancements arising from the different values of $K$, shown for each dataset separately, and reported with both AUC and B-ACC metrics. Each data point represents the average result across all three forward LLMs along with all their corresponding backward LLMs (i.e. the average of each column in tables 3,9 and 10). As can be seen, the data reveals a consistent trend wherein the cumulative improvements exhibit a proportional relationship with the size of $K$.
@@
We extend our investigation to varying temperatures for the backward process. For each index $i\in\text{range}(K)$, the InterrogateLLM method utilizes a variable temperature $T_{i}$ as defined in Eq.(8). This temperature adjustment aimed to augment the creativity and stochastic aspects of the backward model $B_{LLM}$ throughout the query reconstruction process, fostering the generation of a more diverse set of reconstructed queries. Tab. 4, and Tab. 6, 7 from the supplementary Sec.A.3, present the results of InterrogateLLM with $K=5$, when using the same temperature through all the backward processes versus using variable temperatures, as proposed in InterrogateLLM. As can be seen, the variable temperature improves the results across most experiments in the Movies datasets, while yielding on-par performance in the Books and GCI datasets (see Tab.6, 7). We hypothesize that the introduction of variable temperature, generating reconstructions with diverse levels of creativity, can be particularly helpful in mitigating instances of mode "collapse", in which certain backward models consistently generate identical reconstructions. In such cases, the incorporation of variable temperature becomes more important. The proposed method utilizes a diverse range of reconstructions. When the vast majority of these diverse reconstructions closely align with the original query, it signifies robust backward processes, better reflecting a non-hallucinated answer and consequently leading to improved accuracy scores. More ablative experiments related to the choices made in Eq. 5 can be found in the sup. Sec.A.1.
6 Conclusion
As a future work, we would like to extend the method to Retrieval Augmented Generation settings, where a query is provided with a retrieved context, and the task is to generate an answer based on the provided information in the context.
@@ A.1 The Average-Max Analysis
For each query, we obtain a list of size K containing cosine similarity scores, representing the similarity between the generated query and the original query. To ascertain the closeness of the generated query to the original, we employ two strategies: the maximum (Max) and the average (Average) of the K cosine similarity scores. Notably, in the majority of cases, the average outperformed the maximum, indicating a more robust measure for query similarity. We attribute this observation to the sensitivity of the maximum strategy to noisy cosine similarity scores, which tend to be prevalent in certain instances. Our analysis suggests that the average is a more reliable indicator of query similarity, providing a stable and effective metric for the evaluation of our reconstruction approach. Full results with avg/max are presented in tables 5, 6, 7, 8, 9 and 10.